# Monocular 3D lane detection for Autonomous Driving: Recent Achievements, Challenges, and Outlooks

### v1

@@ V-B4 Adverse Weather Conditions
3D-LaneNet [24] 53 RTX 2080 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [25] 50 RTX 2080 32.3 25.4 33.5 28.1 18.7 21.4 31.0
PersFormer [12] 21 Tesla V100 50.5 42.4 55.6 48.6 46.6 40.0 50.7
CurveFormer [7] - - 50.5 45.2 56.6 49.7 49.1 42.9 45.4
Anchor3DLane [31] 87.3 RTX 2080Ti 54.3 47.2 58.0 52.7 48.7 45.8 51.7
BEV-LaneDet [90] 102 Tesla V100 58.4 48.7 63.1 53.4 53.4 50.3 53.7
Efficient Transformer [15] 72 Tesla V100 63.8 57.6 73.2 57.3 59.7 57.0 63.9
GroupLane [48] 38.2 RTX 2080 64.1 - - - - - -
LATR [53] 17.75 Tesla A100 61.9 55.2 68.2 57.1 55.4 52.3 61.5
DecoupleLane [26] - - 51.2 43.5 57.3 - 48.9 43.5 -
CurveFormer++ [8] 11.4 RTX 3090 52.7 48.3 59.4 50.6 48.4 45.0 48.1
@@
3D-LaneNet [24] 53 RTX 2080 44.73 61.46 35.16 0.127
Gen-LaneNet [25] 60 RTX 2080 45.59 63.95 35.42 0.121
SALAD [107] - - 64.07 75.90 55.42 0.098
PersFormer [12] 21 Tesla V100 74.33 80.30 69.18 0.074
WS-3D-Lane [3] 26.7 RTX 3080Ti 77.02 84.51 70.75 0.058
Anchor3DLane [31] 87.3 RTX 2080Ti 74.87 80.85 69.71 0.060
Efficient Transformer [15] 72 Tesla V100 80.84 84.50 77.48 0.056
GroupLane [48] 38.2 RTX 2080 80.73 82.56 78.90 0.053
LATR [53] 17.75 Tesla A100 80.59 86.12 75.73 0.052
DecoupleLane [26] - - 75.07 81.19 69.26 0.062
CurveFormer++ [8] 11.4 RTX 3090 77.85 81.06 74.89 0.084
@@ V-B6 Towards More Accurate and Efficient 3D Lane Detection Methods for Autonomous Driving
The qualitative performance of current 3D lane detection techniques is shown in Table IV. We can observe that only a few methods manage to strike a balance between model accuracy and inference latency. The experimental results of these methods indicate the need for further improvements to alleviate computational burdens while maintaining their unparalleled performance. Additionally, We added noise to the dataset to simulate adverse weather conditions, and tested the performance of the 3D lane line detection algorithm on these challenging datasets, the performance of the algorithm was not satisfactory, as depicted in Figure 5. We found that these models should also be evaluated for knowledge transferability and generalization across different datasets. Moreover, the reported time complexities in Table [?] suggest that some methods can achieve real-time execution when deployed on GPU devices. Nevertheless, considering the constrained computational resources in today’s autonomous driving telemetry, the focus of 3D lane detection methods should also shift towards computational complexity.
@@
RGB cameras are limited by their imaging principles, resulting in poor image quality in high-speed or low-light scenes. Fortunately, event cameras can overcome this limitation. Event cameras are visual sensors with high temporal resolution, high dynamic range, low latency, and low energy consumption [99]. Unlike traditional cameras that capture images based on the intensity and color of light, event cameras capture images based on changes in light intensity [22, 23]. Therefore, event cameras can capture images in low-light scenes as long as there are changes in light intensity. As a result, there is significant and extensive research potential in the field of 3D lane detection based on event cameras, including the development of datasets for 3D lane detection with event cameras and Algorithms specifically designed for 3D lane detection using event cameras or a combination of event and RGB cameras.
@@ V-B8 Uncertainty-aware 3D Lane Detection
Over the past few years, DNNs have garnered significant success in numerous computer vision assignments, solidifying their status as indispensable instruments for proficient automated perception. Despite consistently delivering exceptional results across different benchmarks and tasks, there remain several significant hurdles to overcome before widespread implementation. Among the most common and well-known critiques of DNNs is their susceptibility to unreliable performance when confronted with shifts in data distribution levels, highlighting the urgent need to rectify this limitation [20]. Currently, the majority of deep learning models provide deterministic outputs, giving a single result. However, in real-world driving scenarios, it is desirable for models to provide uncertainty estimates for their predictions [62, 38, 57]. Downstream decision-making modules can then use this uncertainty information to make more reasonable and safe driving commands. For example, in the context of 3D lane detection, if the model outputs a lane position with high uncertainty, we should approach the model’s detection result with skepticism, and a conservative driving style should be adopted. Conversely, if the model’s output has low uncertainty, we can have confidence in the algorithm’s prediction and make more confident driving decisions.

### v2

@@ V-B4 Adverse Weather Conditions
3D-LaneNet [garnett20193d] 53 RTX 2080 44.1 40.8 46.5 47.5 41.5 32.1 41.7
GenLaneNet [guo2020gen] 50 RTX 2080 32.3 25.4 33.5 28.1 18.7 21.4 31.0
PersFormer [chen2022persformer] 21 Tesla V100 50.5 42.4 55.6 48.6 46.6 40.0 50.7
CurveFormer [bai2022curveformer] - - 50.5 45.2 56.6 49.7 49.1 42.9 45.4
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 54.3 47.2 58.0 52.7 48.7 45.8 51.7
BEV-LaneDet [wang2023bevlanedet] 102 Tesla V100 58.4 48.7 63.1 53.4 53.4 50.3 53.7
Efficient Transformer [chen2023efficient] 72 Tesla V100 63.8 57.6 73.2 57.3 59.7 57.0 63.9
GroupLane [li2023grouplane] 38.2 RTX 2080 64.1 - - - - - -
LATR [luo2023latr] 17.75 Tesla A100 61.9 55.2 68.2 57.1 55.4 52.3 61.5
DecoupleLane [han2023decoupling] - - 51.2 43.5 57.3 - 48.9 43.5 -
CurveFormer++ [bai2024curveformer++] 11.4 RTX 3090 52.7 48.3 59.4 50.6 48.4 45.0 48.1
@@
3D-LaneNet [garnett20193d] 53 RTX 2080 44.73 61.46 35.16 0.127
Gen-LaneNet [guo2020gen] 60 RTX 2080 45.59 63.95 35.42 0.121
SALAD [yan2022once] - - 64.07 75.90 55.42 0.098
PersFormer [chen2022persformer] 21 Tesla V100 74.33 80.30 69.18 0.074
WS-3D-Lane [ai2022ws] 26.7 RTX 3080Ti 77.02 84.51 70.75 0.058
Anchor3DLane [huang2023anchor3dlane] 87.3 RTX 2080Ti 74.87 80.85 69.71 0.060
Efficient Transformer [chen2023efficient] 72 Tesla V100 80.84 84.50 77.48 0.056
GroupLane [li2023grouplane] 38.2 RTX 2080 80.73 82.56 78.90 0.053
LATR [luo2023latr] 17.75 Tesla A100 80.59 86.12 75.73 0.052
DecoupleLane [han2023decoupling] - - 75.07 81.19 69.26 0.062
CurveFormer++ [bai2024curveformer++] 11.4 RTX 3090 77.85 81.06 74.89 0.084
@@ V-B6 Towards More Accurate and Efficient 3D Lane Detection Methods for Autonomous Driving
The qualitative performance of current 3D lane detection techniques is shown in Table IV. We can observe that only a few methods manage to strike a balance between model accuracy and inference latency. The experimental results of these methods indicate the need for further improvements to alleviate computational burdens while maintaining their unparalleled performance. Additionally, we selected some challenging data from the 3D lane detection dataset and tested the performance of the 3D lane line detection algorithm on these challenging data samples. However, the performance of the algorithm was not satisfactory,as depicted in Fig. 9. Improving the algorithm’s detection performance in extreme weather conditions is also crucial. Moreover, the reported time complexities in Table IV,V and VI suggest that some methods can achieve real-time execution when deployed on GPU devices. Nevertheless, considering the constrained computational resources in today’s autonomous driving system, the focus of 3D lane detection methods should also shift towards computational complexity.
@@
RGB cameras are limited by their imaging principles, resulting in poor image quality in high-speed or low-light scenes. Fortunately, event cameras can overcome this limitation. Event cameras are visual sensors with high temporal resolution, high dynamic range, low latency, and low energy consumption [enwiki:1202151981]. Unlike traditional cameras that capture images based on the intensity and color of light, event cameras capture images based on changes in light intensity [gallego2020event, gao2022vector]. Therefore, event cameras can capture images in low-light scenes as long as there are changes in light intensity. Currently, there is limited research on 3D lane detection based on event cameras. We believe that there is significant and extensive research potential in the field of 3D lane detection using event cameras, including the development of datasets specifically for 3D lane detection with event cameras and the design of algorithms tailored for 3D lane detection using event cameras alone or in combination with RGB cameras.
@@ V-B8 Uncertainty-aware 3D Lane Detection
Over the past few years, DNNs have garnered significant success in numerous computer vision assignments, solidifying their status as indispensable instruments for proficient automated perception. Despite consistently delivering exceptional results across different benchmarks and tasks, There are still several significant hurdles that need to be overcome before widespread implementation. Among the most common and well-known critiques of DNNs is their susceptibility to unreliable performance when confronted with shifts in data distribution levels, highlighting the urgent need to rectify this limitation [franchi2022muad]. Currently, the majority of deep learning models provide deterministic outputs, giving a single result. However, in real-world driving scenarios, it is desirable for models to provide uncertainty estimates for their predictions [mcallister2017concrete, kendall2017uncertainties, ma2023self]. Downstream decision-making modules can then use this uncertainty information to make more reasonable and safe driving commands. For example, in the context of 3D lane detection, if the model outputs a lane position with high uncertainty, we should approach the model’s detection result with skepticism, and a conservative driving style should be adopted. Conversely, if the model’s output has low uncertainty, we can have confidence in the algorithm’s prediction and make more confident driving decisions.